OpenStudy (anonymous):

A mixture of hydrogen and oxygen are in the volume ratio 1:4 the initial composition of H2 to O2 in the mixture Diffusing initially is?

OpenStudy (anonymous):

@ghazi @Kryten @sauravshakya

OpenStudy (ghazi):

\[\frac{ V _{H} }{ V _{O} }=\frac{ 1 }{ 4 }\]

OpenStudy (anonymous):

so it is: n(H2):n(O2)=1:4 and since those are gasses you can say that: V(% H2)= 1/5 V(% O2)= 4/5

OpenStudy (ghazi):

@Kryten it is in volume ratio and may be there is a container of that volume which holds gases in the ratio

OpenStudy (anonymous):

volume ratio is same as molar ratio cause we are talking about gasses! so that ratio is basicly saying that on every one part of hydrogen there are 4 parts of oxygen....

OpenStudy (ghazi):

hmm

OpenStudy (anonymous):

A mixture of hydrogen and oxygen are in the volume ratio 1:4 the initial composition of H2 to O2 in the mixture Diffusing initially is?

OpenStudy (anonymous):

i forgot to add Diffusing***

OpenStudy (anonymous):

so you are asking for composition of mixture before diffusing? and it is: 1/5 = 0,2 or 20% H2 and 4/5 = 0,8 or 80% O2